# Universal Anomaly Detection at the LHC: Transforming Optimal Classifiers and the DDD Method

### v2

@@ Acknowledgements
This appendix supplements the main text with additional detailed tables and analysis results. The tables included here summarise the performance metrics for the various machine learning models and techniques discussed in this paper, with a focus on their application to the 4-top channel and the DarkMachines challenge channels. Each table contains data on key performance indicators such as the AUC, $\epsilon_{S}$ and $\epsilon_{B}$.
Appendix A Tables for performance comparisons
A.1 4-top channel
DeepSVDD DDD DROCC
$t\bar{t}t\bar{t}$ AUC 0.606 0.518 0.480 0.754 0.732 0.715 0.550 0.590
$\epsilon_{B}(\epsilon_{S}=0.01)$ $<0.001$ 0.013 0.014 0.003 0.001 0.003 0.004 0.004
$\epsilon_{S}(\epsilon_{B}=0.01)$ 0.124 0.007 0.007 0.033 0.043 0.025 0.022 0.026
Table 5: In the 4-top results, the metrics for the SVDD and DDD methods are calculated based on the average predictions of four ensembled models. Bold numbers represent the highest performance.
A.2 Dark Machines channels

### v3

@@ 6 Conclusions
Furthermore, it is noteworthy that the DDD model not only showed high effectiveness in detecting anomalies in different signals and channels among the evaluated methods, but also ranked among the three best performing models in the DarkMachines challenge.
@@ Acknowledgements
This appendix supplements the main text with additional tables and analysis results. These tables provide a detailed summary of the performance metrics for various machine learning models and techniques applied to the DarkMachines challenge channels. Each table includes key metrics such as the area under the ROC curve (AUC), signal efficiency ($\epsilon_{S}$), and background efficiency ($\epsilon_{B}$).
Appendix A Performance Comparison Tables
A.1 Dark Machines channels
Tables 4, 5, 6, and 7 present detailed results for the Dark Machines channels. These include metrics like the AUC and background efficiencies for various signal processes. The information in these tables is used to evaluate and compare the performance of the methods and architectures discussed in the main text.